If the length and breadth of a room are increased by 1 metre, the area is increased by 21 metres. If the length is increased by 1 metre , the area is dcreased by 5 square metres . Find the perimeter of the room.
Answer by mathstutor494(120)

Let the length and breadth of a room be x & y
Area of the room = x * y
When the length and breadth of a room are increased by 1 metre, then the new dimensions of the room would be
Length : x+1
Width : y+1
New area = (x+1)(y+1)
Since the area is increased by 21 metres, therefore
(x+1)(y+1)- xy = 21...........(1)
Further when the length is increased by 1 metre, then the area would be (x+1) * y
Now when the length is increased by 1 metre , the area is dcreased by 5 square metres
(x+1) * y -xy = 21-5............(2)
Simplifying & solving eqns 1 & 2 results in
x=4 and y=16
Perimeter of the room = 2*(Length+width)
= 2*(4+16)
Ans : 40 meter